A soda manufacturer is interested in determining wheter its bottling machine tends to overfill. Each bottle is supposed to contain 12 ounces of fluid. A random sample of 25 bottles was taken and found that the mean amount of soda of the sample of bottles is 12.2 ounces with a standard deviation of 0.4 ounces. If the manufacturer decides on a significance level of 0.05 test, should the null hypothesis (u=12 ounces) be rejected?





1
Expert's answer
2022-06-06T08:52:03-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=12"

"H_1:\mu\not=12"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.05," "df=n-1=24" and the critical value for a two-tailed test is "t_c = 2.063899."

The rejection region for this two-tailed test is "R = \{t:|t|>2.063899\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{12.2-12}{0.4/\sqrt{25}}=2.5"


Since it is observed that "|t|=2.5> 2.063899=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for two-tailed, "df=24" degrees of freedom, "t=2.5" is "p= 0.019654," and since "p= 0.019654<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is different than 12, at the "\alpha = 0.05" significance level.